The server that hosts the iris.eecs.berkeley.edu/network and iris.eecs.berkeley.edu/roster applications is down for maintenance on Monday, May 18th from 5pm-6pm.
The EECS Network is currently disconnected from the rest of the Internet. The EECS Network team is currently investigating.
The Soda Hall machine room has lost power at approximately 11:50am on Dec 20, 2017, during what was supposed to be routine work on the room’s UPS. This has caused an interruption in many IRIS services. Staff are bringing services back as soon as they can.
The Soda Hall machine room has lost power at approximately 3:10pm on Dec 1, 2017, during what was supposed to be routine work on the room’s UPS. This has caused an interruption in many IRIS services. Staff are bringing services back as soon as they can.
News/ Campus losing power, IRIS services shutting down
BY ROB MCNICHOLAS
PUBLISHED 2:59 P.M. AUG 02, 2017
Due to a fire threatening our transformers, PG&E has advised campus we will be losing power.
IRIS is taking proactive steps to shut down as much equipment safely before we totally lose power.
See https://www.nixle.us/9HHZX for the warning from UCPD.
A power outage to Soda Hall which began around 12:30pm on June 21, 2017 has taken out many IRIS services. This backup website will be updated until our main website is working again.
Power has been restored to Soda Hall and staff are bringing up network equipment now. The power conditioner in the main Soda machine room is having problems recovering. Electricians are on site assisting.
This incident was fully resolved around 12:30am on June 22. See this IRIS news article for details.
Routine RedHat Linux operating system patches have been released which require system reboots. Starting at 6pm Wednesday, April 12, IDSG will be patching and rebooting affected servers. The following services will experience brief outages of 10 minutes or less between 6pm and 6:30pm.
- Department mailing list services (lists.eecs.berkeley.edu)
- Department jabber/xmpp service (jabber.eecs.berkeley.edu)
- IRIS website (iris.eecs.berkeley.edu)
The EECS networks became unreachable around 1:20pm on Monday, April 10. Staff are investigating.
Update (2:15pm): Staff investigation found the problem to be at or with one of the department border firewall devices. The secondary firewall device has been made primary now, and traffic is routing properly. Staff will further examine the problem firewall device, to narrow down the specific cause.
IRIS will be migrating our help request tracking system to a new server tonight, Monday March 27 between 6pm and 8pm. Some parts of the IRIS website will be unavailable during this time.
Originally Published 11:56 a.m. Oct 26, 2016 on IRIS News
Services Affected: Home Directory Storage, IRIS Website, Project Storage, Repo Service, and Unix Login Server
Begins 10:00 a.m. Nov 24, 2016
On the morning of Thursday, November 24 (Thanksgiving Day), we will be performing some disruptive maintenance on the department fileserver:
- Recable to new network datacenter switches.
- Upgrade cluster switches firmware.
- Upgrade NetApp OS.
Department fileservice will be unavailable during the maintenance window, approximately 10am until 1pm. Many services that rely on access to department fileservice (home directories, webservice, repo service, etc), will either be offline or will fail to work during this window. All NFS and CIFS mounts will be affected.
We expect most, if not all, clients to recover automatically once the fileservers are back online. This news item will be updated to indicate when the maintenance is completed.
Systems still experiencing mount issues after the maintenance window, would need to remount or reboot.
FILESERVER MAINTENANCE COMPLETED
Pathma Venasithamby at 12:37 p.m. Nov 24, 2016
Scheduled Maintenance Resolved
Systems on new datacenter network switches. All firmware updates completed. Netapp OS upgraded on all nodes. If there are any NFS or CIFS issues please reconnect and followup with helpdesk if there are outstanding issues.